Three capacitors of capacitance 2µF, 4µF and 6µF respectively are connected in series to a 220V dc supply. Find total capacitance of the combination.
1
Expert's answer
2018-11-12T15:54:09-0500

We can find the total capacitance of the combination from the formula:

1/C_tot =1/C_1 +1/C_2 +1/C_3 ,

C_tot=1/(1/C_1 +1/C_2 +1/C_3 )=1/(1/(2 μF)+1/(4 μF)+1/(6 μF))=1.1 μF.